OpenStudy (anonymous):

The problem above shows some values for the function f. If f(x)=ka to the x power for some constants k and a, what is the value of a?

OpenStudy (anonymous):

x= -1,0,1. f(x)=1/8,1/2,2

OpenStudy (anonymous):

tips?

OpenStudy (chihiroasleaf):

f(x) = ka to the x power, so \[f(x) = \left( ka \right)^{x}\] or \[f(x)=ka ^{x}\] you just need to substitute the value of x into the function, then you'll have three equations, solve the equation to find k and a

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

1/4 equal a when k=1/32?

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

a=4 when k=1/8?

OpenStudy (anonymous):

that's the correct answer in the book

OpenStudy (anonymous):

but why?

OpenStudy (chihiroasleaf):

are you sure that it's the result? the given informations are f(-1) = 1/8 ; f(0)=1/2 ; f(1) = 2, right?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

don't you just plug in values for f(x)=ka to the x power?

OpenStudy (anonymous):

1/8=k4 to the negative one power

OpenStudy (anonymous):

1/8=1/4k

OpenStudy (anonymous):

k=1/2

OpenStudy (anonymous):

1/8=1/2 times 1/4

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

are did I do it incorrectly?

OpenStudy (chihiroasleaf):

where do you get 1/8 = k4 and 1/8 = 1/4k ?

OpenStudy (anonymous):

let me try out a simpler equatio

OpenStudy (anonymous):

2=k4 to the first power

OpenStudy (anonymous):

2=4k

OpenStudy (anonymous):

1/2=k

OpenStudy (anonymous):

2=1/2 times 4=2

OpenStudy (chihiroasleaf):

\[f(x)= ka ^{x}\] \[f(-1)=\frac{ 1 }{ 8 }\] \[ka^{-1}=\frac{1}{8}\] \[frac{k}{8}=\frac{1}{8}\]

OpenStudy (anonymous):

wah?

OpenStudy (anonymous):

I thought i Just had to plug in values for the equation f(x)=ka to the x power?

OpenStudy (anonymous):

doesn't f(x) equal y?

OpenStudy (chihiroasleaf):

*the last row \[\frac{ k }{ a }=\frac{ 1 }{8 }\]

OpenStudy (anonymous):

why doesn't 2= k(4 to the first power) work?

OpenStudy (chihiroasleaf):

yes, f(x) equal to y you want to find a right?

OpenStudy (anonymous):

yeah....

OpenStudy (anonymous):

2=4k

OpenStudy (anonymous):

1/2=k, right?

OpenStudy (chihiroasleaf):

yes, k = 1/2 but I get it by using the second information f(0) = 1/2, so k x a^0 = 1/2 k = 1/2

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

do

OpenStudy (anonymous):

so

OpenStudy (anonymous):

1/2=ax to the 0 power

OpenStudy (anonymous):

1/2=a

OpenStudy (anonymous):

k to the 0 power

OpenStudy (anonymous):

the answer is D

OpenStudy (anonymous):

4

OpenStudy (anonymous):

not 1/2...

OpenStudy (anonymous):

wait a sec...

OpenStudy (anonymous):

1/2=ka to the 0 power

OpenStudy (anonymous):

1/2=k

OpenStudy (anonymous):

1/2=1/2a to the x power...

OpenStudy (chihiroasleaf):

yes, then plug in the value of k into the first equation you have, k/a = 1/8 , then you'll get a

OpenStudy (anonymous):

you can do that?

OpenStudy (anonymous):

could you do 1/2=1/2(4 to the 0 power)

OpenStudy (anonymous):

could you do that?

OpenStudy (anonymous):

I got k=1/2

OpenStudy (anonymous):

tips?

OpenStudy (chihiroasleaf):

if you use that equation, any value of a will satisfy the equation since a^0 = 1, so, you cannot use that equation to find a.., yes, k = 1/2, then to find a , plug in k = 1/2 into k/a = 1/8

OpenStudy (anonymous):

where did you get a, though?

OpenStudy (anonymous):

where did you get the 1/8?

OpenStudy (chihiroasleaf):

you have f(-1) = 1/8 so, k x a^(-1) = 1/8 k/a = 1/8

OpenStudy (anonymous):

so a=4?

OpenStudy (anonymous):

I think I understand your reasoning now

OpenStudy (anonymous):

ty

OpenStudy (chihiroasleaf):

yes

OpenStudy (anonymous):

next problem?